The pH of a solution is 4.62 plusminus 0.03. What
Answer #1

part a:
charge =-345 e=-345*1.6*10^(-19) coloumbs

=5.52*10^(-17) C

length of the arc=2*pi*radius*angle/(360 degrees)

=2*pi*0.07*73/360=0.089186 m

then linear charge density=total charge/total length

=-5.52*10^(-17)/0.089186

=-6.1893*10^(-16) C/m


part b:

surface charge density=charge/surface area

=-5.52*10^(-17)/(pi*radius^2)

=-5.52*10^(-17)/(pi*0.022^2)=-3.6303*10^(-14) C/m^2

part c:

surface charge density=charge/surface area

=-5.52*10^(-17)/(4*pi*radius^2)

=-5.52*10^(-17)/(4*pi*0.041^2)

=-2.6131*10^(-15) C/m^2

part d:

volume charge density=charge/volume

=-5.52*10^(-17)/((4/3)*pi*radius^3)

=-5.52*10^(-17)/((4/3)*pi*0.026^3)

=-7.4977*10^(-13) C/m^3